# EmoDynamiX: Emotional Support Dialogue Strategy Prediction by Modelling MiXed Emotions and Discourse Dynamics

### v2

@@ Abstract
Designing emotionally intelligent conversational systems to provide comfort and advice to people experiencing distress is a compelling area of research. Recently, with advancements in large language models (LLMs), end-to-end dialogue agents without explicit strategy prediction steps have become prevalent. However, implicit strategy planning lacks transparency, and recent studies show that LLMs’ inherent preference bias towards certain socio-emotional strategies hinders the delivery of high-quality emotional support. To address this challenge, we propose decoupling strategy prediction from language generation, and introduce a novel dialogue strategy prediction framework, EmoDynamiX, which models the discourse dynamics between user fine-grained emotions and system strategies using a heterogeneous graph for better performance and transparency [^1] Our code is available at https://anonymous.4open.science/r/EmoDynamiX-v2-F6C6.. Experimental results on two ESC datasets show EmoDynamiX outperforms previous state-of-the-art methods with a significant margin (better proficiency and lower preference bias). Our approach also exhibits better transparency by allowing backtracing of decision making.
EmoDynamiX: Emotional Support Dialogue Strategy Prediction by Modelling MiXed Emotions and Discourse Dynamics
Chenwei Wan1        Matthieu Labeau1        Chloé Clavel2 1LTCI, Télécom Paris, Institut Polytechnique de Paris     2Inria Paris 1{chenwei.wan, matthieu.labeau}@telecom-paris.fr 2chloe.clavel@inria.fr
@@ 1 Introduction
Figure 1: This figure shows a possible modular ESC dialogue system, in which our dialogue strategy prediction framework outputs the next dialogue strategy to be used to guide an external generative model.
@@
However, implicit dialogue strategy planning with LLMs faces two challenges. First, transparency is often lacking in such "black-box" decision-making processes. Second, recent studies show that preference biases inherited from pre-training data often cause LLMs to struggle with balancing social-oriented and task-oriented goals. Abulimiti et al. (2023) found that in peer-tutoring dialogues, LLMs like ChatGPT frequently prioritize non-hedging strategies, even in situations where hedging strategies would be more appropriate for repairing low rapport between peers. Similarly, Kang et al. (2024) observed that LLMs’ strong predisposition towards certain strategies can undermine the outcome of the current stage of ESC. As a result, the overall objectives may be significantly compromised.
To address this limitation, introducing external strategy planners, which offer greater controllability by enabling us to explicitly exclude inappropriate strategies in specific contexts, stands out to be a promising solution. It has been evidenced by both automatic metrics and human evaluations that, an external decision agent can more effectively mitigate preference bias and improve emotional support quality compared to relying solely on enhancing end-to-end LLMs (Kang et al., 2024). Therefore, we treat socio-emotional strategy prediction as an independent task, as previously explored by Vanel et al. (2023). We illustrate the scope of this task in Figure 1. From this perspective, we raise the following research questions:
  • RQ1: Can we better enhance performance and transparency in socio-emotional dialogue strategy prediction of modular systems by smartly designing dedicated frameworks, rather than prompting or fine-tuning LLMs?
  • RQ2: Given the importance of emotional intelligence in delivering effective emotional support, can we boost strategy prediction in ESC by integrating the task of Emotion Recognition in Conversations (ERC)?
In addressing RQ1, we introduce EmoDynamiX, a decision agent that integrates multiple expert models and incorporates a heterogeneous graph learning module to capture the dynamic interactions between system strategies and user emotions. With graphs, we backtrace the decision-making process, making a step towards greater transparency. We also utilize dummy nodes (Liu et al., 2022; Scarselli et al., 2008) for role-aware information aggregation, enhancing the overall performance.
@@ 4.2 Heterogeneous Graph Learning
Edge Types: The edges in our heterogeneous graph fulfill dual roles: they model the discourse dependencies between dialogue turns and facilitate information aggregation towards the dummy node. Discourse dependencies correspond to edges between and within emotion and system strategy nodes: we follow Asher et al. (2016)’s definition (which include categories such as Comment and Elaboration) and pre-train a discourse parser as proposed by Chi and Rudnicky (2022) on the multi-party discourse dataset STAC (Asher et al., 2016). Details regarding the training procedure can be found in Appendix G. We note $\mathcal{R}_{\text{Discourse}}$ the set of possible dependencies; then, $\forall(i,j)$ such that $1\leq i,j\leq N-1,\langle v_{i},v_{j}\rangle\in\mathcal{R}_{\text{Discourse}}$. We give more details on these dependencies in Appendix A.2. The remaining edges are aggregating information from system strategy nodes, which we call self-reference: $\forall i\in I_{\texttt{agent}},\langle v_{i},v_{N}\rangle=r_{\text{self}}$; and from user emotion nodes, which we call inter-reference: $\forall i\in I_{\texttt{user}},\langle v_{i},v_{N}\rangle=r_{\text{inter}}$. We note $\mathcal{R}=\mathcal{R}_{\text{Discourse}}\cup\{r_{\text{self}},r_{\text{inter}}\}$ the set of edge types.
@@ 4.2.1 User Emotion Node Embedding: Mixed Emotion Method
For DailyDialog, we categorize emotions into seven groups: Ekman’s six basic emotions plus Neutral, collectively referred to as $\mathcal{E}$. Additional statistical details about DailyDialog, along with our motivation for selecting it, are available in Appendix A.2. Detailed information regarding the implementation and training hyperparameters can be found in Appendix G.
@@
To utilize the information in the emotion label distribution more effectively, we employ a learnable temperature parameter $\tau$; details on the impact of the initialization of $\tau$ can be found in Appendix I.
@@ 4.3 Next Dialogue Strategy Prediction
EmoDynamiX ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$
@@
5 Experiments
5.1 Experimental Setups
Datasets To make our model learn strategies beneficial for both social and task-oriented goals, we select two ESC datasets (in English) where dialogues have been human-evaluated and filtered to ensure that the conversational outcomes are positive and the strategies applied are socially appropriate: (i) ESConv (Liu et al., 2021), an ESC dataset annotated by trained crowd-workers. It comprises 1,300 dialogues and features 8 dialogue strategies. We adhere to the official train/dev/test split [^2] https://huggingface.co/datasets/thu-coai/esconv.; (ii) AnnoMI (Wu et al., 2022), an expert-annotated counselling dataset. It includes 133 dialogues and features 9 therapist strategies. We make the train/dev/test split with an 8:1:1 ratio. Detailed dataset statistics are provided in Appendix A.1. For both datasets, we set the context window size to 5 utterances, resulting in 18,376 samples for ESConv and 4,442 samples for AnnoMI.
Baselines To provide extensive comparisons, we choose baselines from three criteria: (i) prompting LLMs SOTA in dialogue tasks, using task description and supplementary information: ChatGPT (OpenAI, 2023) and LLaMA3-70B (Meta, 2024) with 2-shot learning (+2 shot) or emotion labels (+ ERC). We excluded Chain-of-Thought (Wei et al., 2024) prompting because it has already been shown to be ineffective for our task (Kang et al., 2024); (ii) fine-tuning LLMs as general-purpose dialogue strategy predictors: RoBERTa (Liu et al., 2019), BART (Lewis et al., 2020) and LLaMA3-8B (Meta, 2024); (iii) specialized models for emotional support dialogue strategy prediction: MISC (Tu et al., 2022), MultiESC (Cheng et al., 2022), KEMI (Deng et al., 2023) and TransESC (Zhao et al., 2023). For more details see Appendix D.
Evaluation Metrics We use macro F1 score ($\mathcal{M}$-F1) and weighted F1 score ($\mathcal{W}$-F1) as metrics for evaluating the proficiency of strategy prediction models, since the ground truth strategies in the two datasets have been validated by human evaluators. Given the unbalanced nature of ESC datasets, the accuracy score is not an ideal choice as it can unfairly favor models that predominantly predict the majority classes. We also incorporate the preference bias score ($\mathcal{B}$) as defined by Kang et al. (2024) to quantify the extent to which the model favors its preferred strategies over non-preferred ones (implementation details in Appendix E). An ideal dialogue strategy predictor should achieve strong F1 scores while minimizing preference bias.
Implementation Details We implemented our proposed method using PyTorch (Paszke et al., 2019), initializing with the pre-trained weights from RoBERTa and employing the tokenization tools from Huggingface Transformers (Wolf et al., 2020). For optimization, we used the AdamW optimizer (Loshchilov and Hutter, 2019). Detailed hyperparameter settings can be found in Appendix H.
5.2 Overall Performance
@@ 6 In-depth Analysis of EmoDynamiX
We further study the disagreements between the predictions and human strategies using the confusion matrix (Appendix J). Determining the appropriate timing for emotion-related dialogue strategies versus task-oriented ones is particularly challenging. The model frequently predicts Providing Suggestions whereas the human strategies are among the three emotion-related ones: Reflection of feelings, Self-disclosure, and Affirmation and Reassurance. Notably, over 60% of emotion-related strategies were categorized as task-oriented ones, highlighting the difficulty in making choices between these two strategy categories, as also discussed by Galland et al. (2022). We further analyze the disagreement patterns between predicted and human strategies. By looking at their correlations with the primary emotion categories of emotion nodes with the highest attention weights (Figure 4), we find that "Neutral" contributes a larger proportion to these disagreements in general compared to its overall representation in the ERC module’s output distribution (59.22%, as shown in Appendix F). This suggests that strategies are easier to predict when the model can pick up the user emotions expressed in the context.
@@ 7 Conclusions
Intent of Technology We insist that conversational AI should not be developed to replace humans. Therefore, it is crucial to maintain a clear line between AI and humans (Ethique et al., 2024). Given that the training data for conversational AI, including the two datasets we selected, are primarily curated by humans, the AI (especially those trained using end-to-end methods) may exhibit human-like behaviors. For instance, in Figure 1, the system learns to utilize Self-disclosure strategy by expressing feelings of loneliness, which is not consistent with ethical recommendations. We stress that strategies leading to such human-like behaviors should be applied cautiously and potentially restricted in real-world applications to ensure safety. We believe that our next strategy prediction approach, which allows us to explicitly set desired behaviors for AI, can provide better control over conversational AI in the future.
@@ A.2 Datasets for Pre-training Expert Models
Appendix B Definitions of Strategies in ESConv
@@
Appendix C Definitions of Strategies in AnnoMI
@@
Appendix D Baselines
ChatGPT (OpenAI, 2023) is an advanced language model with 175 billion parameters developed by OpenAI. Using reinforcement learning from human feedback (RLHF), it generates human-like text and excels in natural language processing tasks such as conversation and content creation. The prompting is constructed with the template in Figure 6 and corresponding strategy definitions in Appendix B or Appendix C. To facilitate few-shot learning, we built a case library using the training data and extracted two examples from this library for each inference on the test data. We employed example extraction based on similarity scores computed from sentence-BERT (Reimers and Gurevych, 2019) embeddings.
@@ Appendix D Baselines
Appendix E Implementation of the Preference Bias Score
@@
Appendix F Output Statistics of the ERC Module
@@
Appendix G Implementation Details for Submodules
Appendix H Hyperparameter Settings
Figure 7: Analysis on the initialized value of $\tau$.
Appendix I Impact of the Initialization of $\tau$
Appendix J Confusion Matrix of EmoDynamiX on ESConv

### v3

# EmoDynamiX[Uncaptioned image]: Emotional Support Dialogue Strategy Prediction by Modelling MiXed Emotions and Discourse Dynamics

@@ Abstract
Designing emotionally intelligent conversational systems to provide comfort and advice to people experiencing distress is a compelling area of research. Recently, with advancements in large language models (LLMs), end-to-end dialogue agents without explicit strategy prediction steps have become prevalent. However, implicit strategy planning lacks transparency, and recent studies show that LLMs’ inherent preference bias towards certain socio-emotional strategies hinders the delivery of high-quality emotional support. To address this challenge, we propose decoupling strategy prediction from language generation, and introduce a novel dialogue strategy prediction framework, EmoDynamiX, which models the discourse dynamics between user fine-grained emotions and system strategies using a heterogeneous graph for better performance and transparency [^1] Our code is available at https://github.com/cw-wan/EmoDynamiX-v2.. Experimental results on two ESC datasets show EmoDynamiX outperforms previous state-of-the-art methods with a significant margin (better proficiency and lower preference bias). Our approach also exhibits better transparency by allowing backtracing of decision making.
EmoDynamiX[Uncaptioned image]: Emotional Support Dialogue Strategy Prediction
by Modelling MiXed Emotions and Discourse Dynamics
Chenwei Wan1,2thanks: Work done during the first author’s internship at Inria.        Matthieu Labeau1        Chloé Clavel2 1LTCI, Télécom Paris, Institut Polytechnique de Paris, France    2Inria Paris, France 1{chenwei.wan, matthieu.labeau}@telecom-paris.fr 2chloe.clavel@inria.fr
@@ 1 Introduction
Figure 1: This figure shows a possible modular ESC dialogue system, within which our dialogue strategy prediction framework outputs the next dialogue strategy to be used to guide an external generative model.
@@
However, implicit dialogue strategy planning with LLMs faces two challenges. First, transparency is often lacking in such implicit decision-making processes due to the well-known "black-box" property of LLMs (Ludan et al., 2023; Chhun et al., 2024; Lu et al., 2024). Second, recent studies show that preference biases inherited from pre-training data often cause LLMs to struggle with balancing social-oriented and task-oriented goals. Abulimiti et al. (2023) found that in peer-tutoring dialogues, LLMs like ChatGPT frequently prioritize non-hedging strategies, even in situations where hedging strategies would be more appropriate for repairing low rapport between peers. Similarly, Kang et al. (2024) observed that LLMs’ strong predisposition towards certain strategies can undermine the outcome of the current stage of ESC. As a result, the overall objectives may be significantly compromised.
To address this limitation, introducing external strategy planners, which offer greater controllability by enabling us to explicitly exclude inappropriate strategies in specific contexts, stands out to be a promising solution. It has been evidenced by both automatic metrics and human evaluations that, an explicit decision-making module can more effectively mitigate preference bias, and improved proficiency in dialogue strategy actually enhances overall generation quality (Kang et al., 2024). This foundational insight leads us to isolate and focus specifically on dialogue strategy prediction—a previously intermediate step—with three primary goals: (1) better alignment with human strategies, (2) reduced preference bias, and (3) improved transparency.
Additionally, as more powerful LLMs and improved controlled generation techniques continue to emerge (Dathathri et al., 2020), focusing on dialogue strategy prediction offers a more economical pathway to addressing current dialogue system limitations. Explicit dialogue strategy prediction can serve as a flexible, plug-and-play module compatible with state-of-the-art LLMs or as a foundational component for future RL-based methods, where strong alignment with human strategies is typically a critical first step (Deng et al., 2024).
Therefore, we treat socio-emotional strategy prediction as an independent task, as previously explored by Vanel et al. (2023). We illustrate the scope of this task in Figure 1. From this perspective, we raise the following research questions:
  • RQ1: Can we build a dedicated framework for socio-emotional dialogue strategy prediction that is more transparent by design, while outperforming prompting or fine-tuning LLMs in terms of proficiency?
  • RQ2: Given the importance of emotional intelligence in delivering effective emotional support, can we boost strategy prediction in ESC by by accounting the user’s emotion using an ERC (Emotion Recognition in Conversations) module?
In addressing RQ1, we introduce EmoDynamiX, a decision-making framework that integrates multiple expert models and incorporates a heterogeneous graph learning module to capture the dynamic interactions between system strategies and user emotions. With graphs, we backtrace the decision-making process, making a step towards greater transparency. We also utilize dummy nodes (Liu et al., 2022; Scarselli et al., 2008) for role-aware information aggregation, enhancing the overall performance.
@@ 4.2 Heterogeneous Graph Learning
Edge Types: The edges in our heterogeneous graph fulfill dual roles: they model the discourse dependencies between dialogue turns and facilitate information aggregation towards the dummy node. Discourse dependencies correspond to edges between and within emotion and system strategy nodes: we follow Asher et al. (2016)’s definition (which include categories such as Comment and Elaboration) and pre-train a discourse parser as proposed by Chi and Rudnicky (2022) on the multi-party discourse dataset STAC (Asher et al., 2016). Details regarding the training procedure can be found in Appendix D.2. We note $\mathcal{R}_{\text{Discourse}}$ the set of possible dependencies; then, $\forall(i,j)$ such that $1\leq i,j\leq N-1,\langle v_{i},v_{j}\rangle\in\mathcal{R}_{\text{Discourse}}$. We give more details on these dependencies in Appendix A.2. The remaining edges are aggregating information from system strategy nodes, which we call self-reference: $\forall i\in I_{\texttt{agent}},\langle v_{i},v_{N}\rangle=r_{\text{self}}$; and from user emotion nodes, which we call inter-reference: $\forall i\in I_{\texttt{user}},\langle v_{i},v_{N}\rangle=r_{\text{inter}}$. We note $\mathcal{R}=\mathcal{R}_{\text{Discourse}}\cup\{r_{\text{self}},r_{\text{inter}}\}$ the set of edge types.
@@ 4.2.1 User Emotion Node Embedding: Mixed Emotion Method
For DailyDialog, we categorize emotions into seven groups: Ekman’s six basic emotions plus Neutral, collectively referred to as $\mathcal{E}$. Additional statistical details about DailyDialog, along with our motivation for selecting it, are available in Appendix A.2. Detailed information regarding the implementation and training hyperparameters can be found in Appendix D.2.
@@
To utilize the information in the emotion label distribution more effectively, we employ a learnable temperature parameter $\tau$; details on the impact of the initialization of $\tau$ can be found in Appendix E.1.
@@ 4.3 Next Dialogue Strategy Prediction
Ours EmoDynamiX ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$
@@
5 Experiments
5.1 Experimental Setups
Datasets To make our model learn strategies beneficial for both social and task-oriented goals, we select two ESC datasets (in English) where dialogues have been human-evaluated and filtered to ensure that the conversational outcomes are positive and the strategies applied are socially appropriate: (i) ESConv (Liu et al., 2021), an ESC dataset annotated by trained crowd-workers. It comprises 1,300 dialogues and features 8 dialogue strategies. We adhere to the official train/dev/test split [^2] https://huggingface.co/datasets/thu-coai/esconv.; (ii) AnnoMI (Wu et al., 2022), an expert-annotated counselling dataset. It includes 133 dialogues and features 9 therapist strategies. We make the train/dev/test split with an 8:1:1 ratio. Detailed dataset statistics are provided in Appendix A.1. For both datasets, we set the context window size to 5 utterances [^3] We employ a sliding window method to generate training samples: for each supporter utterance in a dialogue, we take the five preceding utterances as the dialogue history and use the current dialogue strategy as the target for prediction., resulting in 18,376 samples for ESConv and 4,442 samples for AnnoMI.
Baselines To provide extensive comparisons, we choose baselines from three criteria: (i) prompting LLMs SOTA in dialogue tasks, using task description and supplementary information: ChatGPT [^4] The version of ChatGPT is gpt-3.5-turbo-0301 (OpenAI, 2023) and LLaMA3-70B (Meta, 2024) with 2-shot learning (+2 shot) or emotion labels (+ ERC). We excluded Chain-of-Thought (Wei et al., 2024) prompting because it has already been shown to be ineffective for our task (Kang et al., 2024); (ii) fine-tuning LLMs as general-purpose dialogue strategy predictors: RoBERTa (Liu et al., 2019), BART (Lewis et al., 2020) and LLaMA3-8B (Meta, 2024); (iii) specialized models for emotional support dialogue strategy prediction: MISC (Tu et al., 2022), MultiESC (Cheng et al., 2022), KEMI (Deng et al., 2023) and TransESC (Zhao et al., 2023). For more details see Appendix C.
Evaluation Metrics We use macro F1 score ($\mathcal{M}$-F1) and weighted F1 score ($\mathcal{W}$-F1) as metrics for evaluating the proficiency of strategy prediction models, since the ground truth strategies in the two datasets have been validated by human evaluators. Given the unbalanced nature of ESC datasets, the accuracy score is not an ideal choice as it can unfairly favor models that predominantly predict the majority classes. We also incorporate the preference bias score ($\mathcal{B}$) as defined by Kang et al. (2024) to quantify the extent to which the model favors its preferred strategies over non-preferred ones (implementation details in Appendix D.1). An ideal dialogue strategy predictor should achieve strong F1 scores while minimizing preference bias.
Implementation Details We implemented our proposed method using PyTorch (Paszke et al., 2019), initializing with the pre-trained weights from RoBERTa and employing the tokenization tools from Huggingface Transformers (Wolf et al., 2020). For optimization, we used the AdamW optimizer (Loshchilov and Hutter, 2019). Detailed hyperparameter settings can be found in Appendix D.3.
5.2 Overall Performance
@@ 6 In-depth Analysis of EmoDynamiX
We further study the disagreements between the predictions and human strategies using the confusion matrix (Appendix E.3). Determining the appropriate timing for emotion-related dialogue strategies versus task-oriented ones is particularly challenging. The model frequently predicts Providing Suggestions whereas the human strategies are among the three emotion-related ones: Reflection of feelings, Self-disclosure, and Affirmation and Reassurance. Notably, over 60% of emotion-related strategies were categorized as task-oriented ones, highlighting the difficulty in making choices between these two strategy categories, as also discussed by Galland et al. (2022). We further analyze the disagreement patterns between predicted and human strategies. By looking at their correlations with the primary emotion categories of emotion nodes with the highest attention weights (Figure 4), we find that "Neutral" contributes a larger proportion to these disagreements in general compared to its overall representation in the ERC module’s output distribution (59.22%, as shown in Appendix E.2). This suggests that strategies are easier to predict when the model can pick up the user emotions expressed in the context.
@@ 7 Conclusions
Intent of Technology We insist that conversational AI should not be developed to replace humans. Therefore, it is crucial to maintain a clear line between AI and humans (Ethique et al., 2024). Given that the training data for conversational AI, including the two datasets we selected, are primarily curated by humans, the AI (especially those trained using end-to-end methods) may exhibit human-like behaviors. For instance, in Figure 1, the system learns to utilize Self-disclosure strategy by expressing feelings of loneliness, which is not consistent with ethical recommendations. We stress that strategies leading to such human-like behaviors should be applied cautiously and potentially restricted in real-world applications to ensure safety. We believe that our approach, which allows us to explicitly set desired behaviors for AI, can provide better control over conversational AI in the future.
@@
Acknowledgement
We thank the anonymous reviewers for their valuable comments. This work was partially funded by the ANR-23-CE23-0033-01 SINNet project.
@@ A.2 Datasets for Pre-training Expert Models
Appendix B Definitions of ESC Strategies
B.1 Strategies in ESConv
@@
B.2 Strategies in AnnoMI
@@
Appendix C Baselines
ChatGPT (OpenAI, 2023) is an advanced language model with 175 billion parameters developed by OpenAI. Using reinforcement learning from human feedback (RLHF), it generates human-like text and excels in natural language processing tasks such as conversation and content creation. The prompting is constructed with the template in Figure 6 and corresponding strategy definitions in Appendix B.1 or Appendix B.2. To facilitate few-shot learning, we built a case library using the training data and extracted two examples from this library for each inference on the test data. We employed example extraction based on similarity scores computed from sentence-BERT (Reimers and Gurevych, 2019) embeddings.
@@ Appendix C Baselines
Appendix D Implementation Details
D.1 Implementation of the Preference Bias Score
@@
D.2 Implementation Details for Submodules
D.3 Hyperparameter Settings
Appendix E Supplementary Materials for Analysis
E.1 Impact of the Initialization of $\tau$
Figure 7: Analysis on the initialized value of $\tau$.
E.2 Output Statistics of the ERC Module
@@
E.3 Confusion Matrix of EmoDynamiX on ESConv